### v1

@@ 7 Discussion and Future work
At the same time, in this section, we list some limitations that V-RECS presents to encourage further research on the topic and opportunities that the availability of V-RECS enables for visualization and visual analytics researchers and practitioners.
7.1 Limitations and challenges
Here we report limitations concerning V-RECS and the methodology it is built on. More specifically we report on:
@@ 7.1 Limitations and challenges
Visualization Literacy Coverage: The fine-tuning of V-RECS and quantitative experimentation are at this moment based on the NCnet dataset, a derivation from the well-known and state-of-the-art NVBench dataset. The choice was made as this is the only dataset capable of supporting the generation and evaluation of the visualization, explanation (E), Captioning (C), and Suggestions for further data exploration (S). At the same time, additional datasets are made available, both in terms of chart collection methods discussed by Chen and Liu in their recent state-of-the-art report on corpora for automating chart generation [2] or the recently available VisText dataset for benchmarking captioning quality [48]. Although, as discussed in Section 5.2, these datasets are not readily usable for our purposes, we foresee the need to create an integrated resource capable of evaluating all the features produced by V-RECS in a combined way, possibly including also datasets like VIS30k [3] and OldOnlineVis [57].
@@
User Evaluation: In its current form, the qualitative evaluation of V-RECS has been conducted through a comparative experiment with 15 expert users. We foresee the need to expand this study to a broader set of participants to explore the subtle differences that could exist in the presentation and significance layers of the EvaLLM evaluation stack and better evaluate the difference in our method between the proprietary GTP model and the open-source fine-tuned model. Given the richness of this activity, we foresee it as a dedicated effort to be conducted in future work.
@@
Form of Explanation: In its current form, the explanation (E) provided to illustrate the generation process is reported only in textual form. As much as it is the standard for LLMs, and V-RECS aligns with it, we see as a limitation the lack of visual support to this explanation or lack of coordination between sub-parts of the explanation and visual elements of the generated chart. In this sense, the exploration of visual methods to support the explanation, as the ones reported by La Rosa et al. [21] or recently by Spinner et al. [43] for LLMs, should be the area for further enhancement of this approach.
7.2 Opportunities
@@
Support for training activities Still referring to the work by Zhao et al. [58], we foresee the opportunity to exploit V-RECS for training activities (referred to as the onboarding stage for Visual Analytics systems) on how to develop a visualization for a specific analysis task, through the support of explanations, or a workflow, exploiting a chain in which at each recommendation stage a new suggestion is selected and given as input for the next recommendation step. This activity would be in line with the recent trend of using LLM for education like the experiment by Chen et al. [6] testing ChatGPT-3.5 and 4 to complete assignments of a visualization course (CS171) and the preliminary work proposed by Joshi et al. [17] on leveraging ChatGPT-3.5 and Bard 2.0 for teaching a visualization technique.
Acknowledgements.
The authors wish to thank the participants who contributed to the manual assessment of V-RECS. This work was in part supported by a grant from Lazio Region, FESR Lazio 2021-2027, project @HOME (# F89J23001050007). This research has also been supported by the AWS Cloud Credit for Research Program, and the OpenAI’s Researcher Access Program.

### v2

@@ VII Conclusions and Future work
At the same time, we list hereafter some limitations that V-RECS presents to encourage further research on the topic and opportunities that the availability of V-RECS enables for visualization and visual analytics researchers and practitioners.
VII-A Limitations and challenges
@@
Visualization Literacy Coverage: The fine-tuning of V-RECS and quantitative experimentation are at this moment based on the NCnet dataset, a derivation from the well-known and state-of-the-art NVBench dataset. As discussed in Section V-B, this choice was made as this is the only dataset capable of supporting the generation and evaluation of the visualization along with its related narratives E,C,S. Although additional datasets have been made available, both in terms of chart collection methods discussed by Chen and Liu in their recent state-of-the-art report on corpora for automating chart generation [52], or the recently available VisText dataset for benchmarking captioning quality [47] these datasets are not readily usable for our purposes, as discussed in Section V-B. We foresee the need for an effort to create an integrated resource capable of evaluating all the features produced by V-RECS in a combined way.
@@
User Evaluation: In its current form, the qualitative evaluation of V-RECS has been conducted through a comparative experiment with 35 annotators. We foresee the need to expand this study to a broader set of participants to explore the subtle differences that could exist in the presentation and significance layers of the EvaLLM evaluation stack and better evaluate the difference in our method between the proprietary GPT model and the open-source fine-tuned model. Given the richness of this activity, we foresee it as a dedicated effort to be conducted in future work.
@@
Form of Narrative: In its current form, the explanation (E) provided to illustrate the generation process is reported only in textual form. As much as it is the standard for LLMs, and V-RECS aligns with it, we see as a limitation the lack of visual support to this explanation or lack of coordination between sub-parts of the explanation and visual elements of the generated chart. In this sense, the exploration of visual methods to support the explanation, as the ones reported by La Rosa et al. [53] or recently by Spinner et al. [54] for LLMs, should be the area for further enhancement of this approach. In addition to that, we note that the answer to a specific query (both the chart and the narrative) is currently unique, while in principle several alternatives could be presented to the user, avoiding the illusion that the output is exclusive, and potentially revealing inconsistencies. The latter is a relevant potential area for further improvement.
VII-B Opportunities
@@
Support for training activities Still referring to the work by Zhao et al. [55], we foresee the opportunity to exploit V-RECS for training activities (referred to as the onboarding stage for Visual Analytics systems) on how to develop a visualization for a specific analysis task, through the support of explanations, or a workflow, exploiting a chain in which at each recommendation stage a new suggestion is selected and given as input for the next recommendation step. This activity would be in line with the recent trend of using LLM for education, like the experiment by Chen et al. [56] testing ChatGPT-3.5 and 4 to complete assignments of a visualization course (CS171) and the preliminary work proposed by Joshi et al. [57] on leveraging ChatGPT-3.5 and Bard 2.0 for teaching a visualization technique.
Acknowledgments
The authors wish to thank the participants who contributed to the manual assessment of V-RECS. This work was in part supported by a grant from Lazio Region, FESR Lazio 2021-2027, project @HOME (# F89J23001050007 CUP B83C23006240002). This research has also been supported by the AWS Cloud Credit for Research Program, and the OpenAI’s Researcher Access Program.